### v2

# Recent Trends of Multimodal Affective Computing: A Survey from an NLP Perspective

Abstract
Multimodal affective computing has garnered increasing attention due to its broad applications in analyzing human behaviors and intentions, especially in text-dominated multimodal affective computing field. This survey presents the recent trends of multimodal affective computing from an NLP perspective through four hot tasks: multimodal sentiment analysis, multimodal emotion recognition in conversation, multimodal aspect-based sentiment analysis and multimodal multi-label emotion recognition. The goal of this survey is to explore the current landscape of multimodal affective research, identify development trends, and highlight the similarities and differences across various tasks, offering a comprehensive report on the recent progress in multimodal affective computing from an NLP perspective. This survey covers the formalization of tasks, provides an overview of relevant works, describes benchmark datasets, and details the evaluation metrics for each task. Additionally, it briefly discusses research in multimodal affective computing involving facial expressions, acoustic signals, physiological signals, and emotion causes. Additionally, we discuss the technical approaches, challenges, and future directions in multimodal affective computing. Finally, we release a repository that compiles related works in multimodal affective computing, providing detailed resources and references for the community [^1] https://github.com/LeMei/Affective-Computing.
I Introduction
Affective computing combines expertise in computer science, psychology, and cognitive science and its goal is to equip machines with the ability to recognize, interpret, and emulate human emotions [1, 2, 3, 4, 5, 6, 7]. Today, the world around us comprises various modalities—we perceive objects visually, hear sounds audibly, feel textures tangibly, smell odors olfactorily, and so forth. A modality refers to the way an experience is perceived or occurs, and is often associated with sensory modalities such as vision or touch, which are essential for communication and sensation. The significant advancements in multimodal learning across various fields [8, 9] garnered increasing attention and accelerated the progress of multimodal affective computing. Multimodal affective computing seeks to develop models capable of interpreting and reasoning sentiment or emotional state over multiple modalities.
In its early stages, researchers of affective computing predominantly focused on unimodal tasks, examining text-based, audio-based, and vision-based affective computing separately. For instance, D-MILN [10] is a textual sentiment classification model, while work [11] utilizes BiLSTM models trained on raw audio to predict the average sentiment of crowd responses. Today, sentiment analysis is widely employed across various modalities for applications such as market research, brand monitoring, customer service analysis, and social media monitoring. Recent advancements in multimedia technology [12, 13, 14, 15] have diversified the channels for information dissemination, with an influx of news, social media platforms like Weibo, and video content. These developments have integrated textual (spoken features), acoustic (rhythm, pitch), and visual (facial attributes) information to comprehensively analyze human emotions. For example, Xu et al. [16] introduced image modality data into traditional text-based aspect-level sentiment analysis, creating the new task of multimodal aspect-based sentiment analysis. Similarly, Wang et al. [17] extended textual emotion-cause pair extraction to a multimodal conversation setting, utilizing multimodal signals to enhance the model’s ability to understand the emotions and their causes.
Multimodal affective computing tasks closely related to several learning paradigms in machine learning, including transfer learning [18, 19, 20], multimodal learning [21, 22], muti-task learning [23, 24, 25] and semantic understanding [26, 27]. Regarding transfer learning, it allows affective analysis models trained in one domain to be adapted for effective performance in different domains. By fine-tuning pre-trained models on limited data from the target domain, these models can be transferred to new domain, thereby enhancing their performance in multimodal affective computing tasks. In multimodal learning, cross-modal attention dynamically aligns and focuses on relevant information from different modalities, enhancing the model’s ability to capture sentiment by highlighting key features and their interactions. In multi-task learning, shared representations across affective computing tasks and modalities improve performance by capturing common sentiment-related features from text, audio, and video. More recently, the studies of multimodal learning have advanced the field by pre-training multimodal models on extensive multimodal datasets over years, further improving the performance on downstream tasks such as multimodal sentiment analysis [28, 29, 30, 31]. With the scaling of the pre-trained model, parameter-efficient transfer learning emerges such as adapter [32], prompt [33], instruction-tuning [34] and in-context learning [35, 36].
More and more works of multimodal affective computing leverage these parameter-efficient transfer learning methods to transfer knowledge from pre-trained models (e.g., unimodal pre-trained model or multimodal pre-trained model) to downstream affective tasks to improve model performance by further fine-tuning the pre-trained model. For instance, Zou et al. [37] design a multimodal prompt Transformer (MPT) to perform cross-modal information fusion. UniMSE [38] proposes an adapter-based modal fusion method, which injects acoustic and visual signals into the T5 model to fuse them with multi-level textual information.
Multimodal affective computing encompasses tasks like sentiment analysis, opinion mining, and emotion recognition using modalities such as text, audio, images, video, physiological signals, and haptic feedback. This survey focuses mainly on three key modalities: natural language, visual signals, and vocal signals. We highlight four main tasks in this survey: Multimodal Sentiment Analysis (MSA), Multimodal Emotion Recognition in Conversation (MERC), Multimodal Aspect-Based Sentiment Analysis (MABSA), and Multimodal Multilabel Emotion Recognition (MMER). A considerable volume of studies exists in the field of multimodal affective computing, and several reviews have been published [39, 40, 41, 15, 42, 43]. However, these reviews primarily focus on specific affective computing tasks or specific single modality and overlook an overview of multimodal affective computing across multiple tasks, and the consistencies and differences among these tasks. The goal of this survey is twofold. First, this survey aims to provide a comprehensive overview of multimodal affective computing for beginners exploring deep learning in emotion analysis, detailing tasks, inputs, outputs, and relevant datasets. Second, it also offers insights for researchers to reflect on past developments, explore future trends, and examine technical approaches, challenges, and research directions in areas such as multimodal sentiment analysis and emotion recognition.
II Organization of this Survey
Section III outlines task formalization and application scenarios for multimodal affective tasks. Section IV introduces feature extraction methods and recent multimodal pre-trained models like (e.g., CLIP, BLIP, BLIP2). Section V analyzes multimodal affective works from the two perspectives: multimodal fusion and multimodal alignment, and shortly summarizes the parameter-efficient transfer methods used for further tuning the pre-trained model. Section VI reviews literature on MSA, MERC, MABSA, and MMER, focusing on multitask learning, pre-trained models, enhanced knowledge, and contextual information. Furthermore, Section VII summarizes multimodal datasets, and Section VIII covers evaluation metrics for each multimodal affective computing task. After the reviews of multimodal affective computing works. Section IX briefly reviews multimodal affective computing works based on facial expressions, acoustic signals, physiological signals, and emotion causes. It also highlights the consistency, differences, and recent trends of multimodal affective computing tasks from an NLP perspective. Section X looks ahead to future work from three aspects of the unification of multimodal affective computing tasks, the incorporation of external knowledge, and affective computing with less-studied modalities. Lastly, Section XI concludes this survey and its contribution to multimodal affective computing community.
III Multimodal Affective Computing Tasks
In this section, we show the definition of each task and discuss their application scenarios. Table I presents basic information, including task input, output, type, and parent task, for each of the four tasks.
III-A Multimodal Sentiment Analysis
Multimodal sentiment analysis (MSA) [44] origins from sentiment analysis (SA) task [45] and it extends SA with the multimodal input. As a key research topics for computers to understand human behaviors, the goal of multimodal sentiment analysis (MSA) is to predict sentiment polarity and sentiment intensity based on multimodal signals [46]. This task belongs to binary classification and regression task.
III-A1 Task Formalization
Given a multimodal signal $I_{i}=\{I^{t}_{i},I^{a}_{i},I^{v}_{i}\}$, we use $I^{m}_{i},m\in\{t,a,v\}$ to represent unimodal raw sequence drawn from the video fragment $i$, where $\{t,a,v\}$ denote the three types of modalities—text, acoustic and visual. Multimodal sentiment analysis aims to predict the real number $y^{r}_{i}\in\mathbb{R}$, where $y^{r}_{i}\in[-3,3]$ reflects the sentiment strength. We feed $I_{i}$ as the model input and train a model to predict $y^{r}_{i}$.
III-A2 Application Scenarios
We categorize multimodal sentiment analysis applications into key areas: social media monitoring, customer feedback, market research, content creation, healthcare, and product reviews. For example, analyzing sentiment in text, images, and videos on social media helps gauge public opinion and monitor brand perception, while analyzing multimedia product reviews can improve personalized recommendations and user satisfaction.
III-B Multimodal Emotion Recognition in Conversation
Initially, MERC [41, 40, 47] extends from emotion recognition in a conversation (ERC) task [48, 49] and it takes multimodal signals as the model inputs instead of signal modality. The goal of MERC is to automatically detect and monitor the emotional states of speakers in a dialogue using multimodal signals like text, audio and vision. In the community of multimodal emotion recognition in a conversation, MERC is a multi-class classification task and then categorizes the given utterance into one basic emotion from a pre-define emotion set.
III-B1 Task Formalization
Given a dialog for $k$ number of utterances, it can be formulated as $U=\{u_{1},\cdots,u_{k}\}$, where $u_{i}=<I_{t}^{i},I_{a}^{i},I_{v}^{i}>$ denotes $i$th utterance of a conversation containing text (text transcript), audio (speech segment) and visual (video clip) modalities, denoted by $\{I_{t}^{i},I_{a}^{i},I_{v}^{i}\}$. We use $Y$ as the label set of $U$ and each utterance can be formalized as follows:
$<U,Y>=\{\{u^{i},y^{i},i\in[1,k]\}\}$ (1)
Here, $y^{i}$ indicates $i$th utterance’s emotion category that is predefined before.
III-B2 Application Scenarios
Multimodal Emotion Recognition in Conversation (MERC) has broad applications across key areas: human-computer interaction, virtual assistants, healthcare, and customer service. (i) In Human-Computer Interaction, MERC enhances user experience by enabling systems to recognize and respond to emotional states, leading to more personalized interactions. (ii) Virtual Assistants and Chatbots benefit from improved emotional understanding, making conversations more natural and engaging. (iii) In Customer Service, MERC helps agents better respond to customer emotions, enhancing satisfaction. Additionally, bio-sensing systems measuring physiological signals like ECG, PPG, EEG, and GSR expand MERC applications in robotics, healthcare, and virtual reality.
III-C Multimodal Aspect-based Sentiment Analysis
Xu et al. [50] are among the first to put forward the new task, aspect based multimodal sentiment analysis. Multimodal aspect-based sentiment analysis (MABSA) is contructed based on aspect-based sentiment analysis in texts [51, 52]. In contrast with MSA and ERC, multimodal aspect-based sentiment analysis performs on fined granularity multimodal signals. MABSA receives texts and vision (image) modalitie as the inputs and outputs the tuple including aspect and its sentiment polarity. This task can be viewed as the classification, tuple extraction and triple extraction tasks. Recently, MABSA has attracted increasing attention. Given an image and corresponding text, MABSA is defined as jointly extracting all aspect terms from image-text pairs and predicting their sentiment polarities, i.e., positive, negative and neutral.
III-C1 Task Formalization
Suppose the multimodal inputs include a textual content $T=\{w_{1},w_{2},...,w_{L}\}$ and an image set $I=\{I_{1},I_{2},\cdots,I_{K}\}$, the goal of MABSA is to predict the sentiment polarities with a given aspect phrase $A=\{a_{1},a_{2},\cdots,a_{N}\}$, where $a_{i}$ denotes the $i$th aspect (e.g., food), $L$ is the length of textual context, $K$ is the number of images, and $N$ is the length of aspect phrase.
III-C2 Application Scenarios
Multimodal aspect-based sentiment analysis (MABSA) focuses on improving products and services by analyzing reviews across text, images, and videos to identify customer opinions on specific aspects. For example, MABSA can assess dining experiences, like food quality or service, to enhance restaurant operations. It also applies to social media, where analyzing mixed content provides deeper insights into public opinion, aiding better decision-making and marketing strategies.
TABLE I: The details of multimodal affective tasks. T,A,V denote text, audio and visual modalities respectively.
Task Input Output Granularity Task type Parent task
MSA T,A,V score,[-3,3] sentence binary classification, regression sentiment analysis [45]
MERC T,A,V one basic emotion utterance multi-class classification emotion recognition [48, 49]
MABSA T,V (aspect, sentiment polarity) aspect classification, tuple extraction, triple extraction aspect-level sentiment analysis [51, 53, 52]
MMER T,A,V two or more basic emotions utterance multi-label multi-class classification emotion recognition [54, 55]
III-D Multimodal Multi-label Emotion Recognition
Multimodal signals may show more than one emotion label, which boosts the rise of a new task: multimodal multi-label emotion recognition (MMER). MMER inherits the characteristic of multimodal emotion recognition and multi-label classification [54, 55]. MMER is developed from multi-label emotion recognition, which predicts two or more basic emotion categories to analysis the given multimodal information and it is a multi-label multi-class classification.
III-D1 Task Formalization
Given a multimodal signal $I_{i}=\{I^{t}_{i},I^{a}_{i},I^{v}_{i}\}$, $I_{i}$ contains three types of modalities—text, audio and visual. Formally, we use $I_{i}^{m}\in\mathbb{R}^{d_{m}\times l_{m}},m\in\{t,a,v\}$ to represent the raw sequence of text, audio, and visual modalities from the sample $i$. $d_{m}$ and $l_{m}$ denote the feature dimension and sequence length of modality $m$. The goal of MMER is to recognize at least one emotion categories from $|L|$ pre-defined label space $Y=\{y_{1},y_{2},\cdots,y_{|L|}\}$ according to the multimodal signal $I_{i}$.
III-D2 Application Scenarios
Multimodal multi-label emotion recognition seeks to create AI systems that can understand and categorize emotions expressed through various modalities simultaneously. This task is challenging due to the complexity and variability of human emotions, differences in emotional expression across individuals and cultures, and the need for effective integration of diverse modalities.
IV Modal Feature Extractor
For multimodal affective computing tasks, the model input typically includes at least two modalities. In this section, we introduce the common feature extractors that transform raw sequences into a feature vectors.
Text Feature Extractor
For text modality, researchers adopt static word embedding methods like Word2Vec [56] and GloVec [57] to initialize word representation. Also, text modality can be encoded into feature vector through pre-trained language models like BERT [58], BART [59], and T5 [60] to extract the text representation. More recently, a collection of foundation language models like LLaMA [61, 62], Mamba [63] emerge and are used for encoding text modality.
Audio Feature Extractor
For audio modality, raw acoustic input needs to be processed into numerical sequential vectors. The common way is to use librosa [^2] https://github.com/librosa/librosa. to extract Mel-spectrogram as audio features. It is the short-term power spectrum of sound and is widely used in modern audio processing. Transformer structure has achieved tremendous success of in the field of NLP and computer vision. Gong et al. [64] propose audio spectrogram Transformer (AST), which converted waveform into a sequence of 128-dimensional log Mel filterbank (fbank) features to encode audio modality.
Vision Feature Extractor
For image modality, researchers can extract fixed T frames from each segment and use effecientNet [65] pre-trained (supervised) on VGGface [^3] https://www.robots.ox.ac.uk/ vgg/software/vgg_face/. and AFEW dataset as vision initial representation. Furthermore, Dosovitskiy et al. [66] propose to use standard Transformer directly to images, which split an image into patches and provide the sequence of linear embeddings of these patches as an input to a Transformer. CLIP [67] jointly trained image and its caption with the contrastive learning, thereby extraction vision features that correspond to texts.
Multimodal Feature Extractor
The emergence of multimodal pre-trained model (MPM) marks a significant advancement in integrating multimodal signals, as demonstrated by groundbreaking developments like GPT-4 [68] and Gemini [69]. Among the open-source innovations, Flamingo [70] represents an early effort to integrate visual features with LLMs using cross-attention layers. BLIP-2 [71] introduces a trainable adaptor module (Q-Former) that efficiently connects a pre-trained image encoder with a pre-trained LLM, ensuring precise alignment of visual and textual information. Similarly, MiniGPT-4 [72] achieves visual and textual alignment through a linear projection layer. InstructBLIP [73] advances the field by focusing on vision-language instruction tuning, building upon BLIP-2, and requiring a deeper understanding and larger datasets for effective training. LLaVA [74] integrates CLIP’s image encoder with LLaMA’s language decoder to enhance instruction tuning capabilities. Akbari et al. [31] train VATT end-to-end from scratch using multimodal contrastive losses and evaluate its performance by the downstream tasks of video action recognition, audio event classification, image classification, and text-to-video retrieval. Based on multimodal pre-trained model, raw modal signals can be used to extract modal features.
V Multimodal Learning on Multimodal Affective Computing
Multimodal learning involves learning representations from different modalities. Generally, the multimodal model should first align the modalities based on their semantics before fusing multimodal signals. After alignment, the model combines multiple modalities into one representation vector.
V-A Preliminary
With the scaling of the pre-trained model, parameter-efficient transfer learning emerges such as adapter [32], prompt [33], instruction-tuning [34] and in-context learning [35, 36]. In this paradigm, instead of adapting pre-trained LMs to downstream tasks via objective engineering, downstream tasks are reformulated to look more like those solved during the original LM training with the help of prompt, instruction-tuning and in-context learning. The use of prompts in Vision Language Models (VLMs) like GPT-4V [68] and Flamingo [70] allows the models to interpret and generate outputs based on combined visual and textual inputs. In contrast with prompt, instruction-tuning belongs to the learning paradigm of prompt. Also, models like InstructBLIP [73] and FLAN [75] have demonstrated that instruction-tuning not only improves the model’s adherence to instructions but also enhances its ability to generalize across tasks. In the community of multimodal affective computing, researchers can leverage these parameter-efficient transfer learning methods (e.g., adapter, prompt and instruction tuning) to transfer knowledge from pre-trained models (e.g., unimodal pre-trained model or multimodal pre-trained model) to downstream affective tasks, further tune the pre-trained model with the affective dataset.
Considering that multimodal affective computing involves multimodal learning, therefore, we analyze multimodal affective computing works from multimodal fusion and multimodal alignment, as shown in Fig. 1.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.8em,font=, where level=2text width=6.3em,font=, where level=3text width=5.5em,font=, [ Multimodal Learning on
Affective Computing, ver [ Multimodal
Fusion (§V-B) [ Cross-modal Learning [ TFN [76], MuLT [77], TCDN [78], CM-BERT [79], HGraph-CL [80], BAFN [81], TeFNA [82],
CMCF-SRNet [83], MultiEMO [84], MM-RBN [85], MAGDRA [86], AMuSE [87]. ,leaf, text width=34.1em ] ] [ Modal Consistency
and Difference [ MMIM [88], MPT [89], MMMIE [90], MISA [91], CoolNet [92], ModalNet [93], MAN [87], TAILOR [94],
AMP [95], STCN [96]. ,leaf, text width=34.1em ] ] [ Multi-stage
Modal Fusion [ TSCL-FHFN [97], HFFN [98], CLMLF [99], RMFN [100], CTFN [101], MCM [102], FmlMSN [103],
ScaleVLAD [104], MUG [105], HFCE [106], MTAG [107], CHFusion [108]. ,leaf, text width=34.1em ] ] ] [ Multimodal
Alignment (§V-C) [ Miss Modality [ MMIN [109], CMAL [110], M2R2 [111], EMMR [112], TFR-Net [113], MRAN [114], VIGAN [115],
TATE [116], IF-MMIN [117], CTFN [101], MTMSA [118], FGR [119], MMTE+AMMTD  [120]. , leaf, text width=34.1em ] ] [ Semantic Alignment [ MuLT [77], ScaleVLAD [104], Robust-MSA [121], HGraph-CL [80], SPIM [122], MA-CMU-SGRNet [123]. ,leaf, text width=34.1em ] ] ] ]
Figure 1: Taxonomy of multimodal affective computing from multimodal fusion and multimodal alignment.
V-B Multimodal Fusion
Multimodal signals are heterogeneous and derived from various information sources, making integrating multimodal signals into one representation essential. Tasi et al. [77] summarize multimodal fusion into early, late or intermediate fusion based on the fusion stage. Early fusion combines features from different modalities at the input level before the model processes them. Late fusion processes features from different modalities separately through individual sub-networks, and the outputs of these sub-networks are combined at a later stage, typically just before making the final decision. Late fusion uses unimodal decision values and combines them using mechanisms such as averaging [124], voting schemes [125], weighting based on channel noise [126] and signal variance [127], or a learned model [128, 6]. The two fusion strategies face some problems. For example, early fusion at the feature level can underrate intra-modal dynamics after the fusion operation, while late fusion at the decision level may struggle to capture inter-modal dynamics before the fusion operation. Different from the previous two methods by combining features from different modalities at intermediate layers of the model learner, Intermediate fusion allows for more interaction between the modalities at different processing stages, potentially leading to richer representations [129, 130, 38]. Based on these fusion strategies, we review multimodal fusion from three aspects: cross-modality learning, modal consistency and difference, and multi-stage modal fusion. Fig. 2 illustrates the three aspects of modal fusion.
V-B1 Cross-modality Learning
Cross-modality learning focuses on the incorporation of inter-modality dependencies and interactions for better modal fusion in representation learning. Early works of multimodal fusion [76] mainly operate geometric manipulation in the feature spaces to fuse multiple modalities. The recent common way of cross-modality learning is to introduce attention-based learning method to model inter-modality and intra-modality interactions. For example, MuLT [77] proposes multimodal Transformer to learn inter-modal interaction. Chen et al. [78] augment the inter-intra modal features with trimodal collaborative interaction and unifies the characteristics of the three modals (inter-modal). Yang et al. [79] propose the cross-modal BERT (CM-BERT), aiming to model the interaction of text and audio modality based on pre-trained BERT model. Lin et al. [80] explore the intricate relations of intra- and inter-modal representations for sentiment extraction. More recently, Tang et al. [81] propose the multimodal dynamic enhanced block to capture the intra-modality sentiment context, which decrease the intra-modality redundancy of auxiliary modalities. Huang et al. [82] propose a Text-centered fusion network with cross-modal attention (TeFNA), a multimodal fusion network that uses crossmodal attention to model unaligned multimodal timing information. In the community of emotion recognition, CMCF-SRNet [83] is a cross-modality context fusion and semantic refinement network, which contains a cross-modal locality-constrained transformer and a graph-based semantic refinement transformer, aiming to explore the multimodal interaction and dependencies among utterances. Shi et al. [84] propose an attention-based correlation-aware multimodal fusion framework MultiEMO, which captures cross-modal mapping relationships across textual, audio and visual modalities based on bidirectional multi-head cross attention layers. In summary, cross-modality learning mainly focuses on modeling the relation between modalities.
Figure 2: Illustration of multimodal fusion from following aspects: 1) cross-modality modal fusion, 2) modal fusion based on modal consistency and difference and 3) multi-stage modal fusion.
V-B2 Modal Consistency and Difference
Modal consistency refers to the shared feature space across different modalities for the same sample, while modal difference highlights the unique information each modality provides. Most multimodal fusion approaches separate representations into modal-invariant (consistency) and modal-specific (difference) components. Modal consistency helps handle missing modalities, while modal difference leverages complementary information from each modality to improve overall data understanding. For example, several works [89, 90] have explored learning modal consistency and difference using contrastive learning. Han et al. [88] maximized the mutual information between modalities and between each modality to explore the modal consistency. Another study [89] proposes a hybrid contrastive learning framework that performs intra-/inter-modal contrastive learning and semi-contrastive learning simultaneously, models cross-modal interactions, preserves inter-class relationships, and reduces the modality gap. Additionally, Zheng et al. [90] combined mutual information maximization between modal pairs with mutual information minimization between input data and corresponding features. This method aims to extract modal-invariant and task-related information. Modal consistency can also be viewed as the process of projecting multiple modalities into a common latent space (modality-invariant representation), while modal difference refers to projecting modalities into modality-specific representation spaces. For example, Hazarika et al. [91] propose a method that projects each modality into both a modality-invariant and a modality-specific space. They implemented a decoder to reconstruct the original modal representation using both modality-invariant and modality-specific features. AMuSE [87] proposes a multimodal attention network to capture cross-modal interactions at various levels of spatial abstraction by jointly learning its interactive bunch of mode-specific peripheral and central networks. For the fine-grain sentiment analysis, Xiao et al. [92] present CoolNet to boost the performance of visual-language models in seamlessly integrating vision and language information. Zhang et al. [93] propose an aspect-level sentiment classification model by exploring modal consistency with fusion discriminant attention network.
Figure 3: Illustration multimodal alignment:(a) semantic alignment and (b) alignment with missing modal fragments.
V-B3 Multi-stage Modal Fusion
Multi-stage multimodal fusion [131, 132] refers to combine modal information extracted from multiple stages or multiple scales to fuse modal representation. Li et al. [97] design a two-stage contrastive learning task, which learns similar features for data with the same emotion category and learns distinguishable features for data with different emotion categories. HFFN [98] divides the procession of multimodal fusion into divide, conquer and combine, which learns local interactions at each local chunk and explores global interactions by conveying information across local interactions. Different from the work of HFFN, Li et al. [99] align and fused the token-level features of text and image and designed label based contrastive learning and data based contrastive learning to capture common features related to sentiment in multimodal data. There are some work [100] decomposed the fusion procession into multiple stages, each of them focused on a subset of multimodal signals for specialized, effective fusion. Also, CTFN [108] presents a novel feature fusion strategy that proceeds in a hierarchical fashion, first fusing the modalities two in two and only then fusing all three modalities. Moreover, the modal fusion at multiple levels has made progress, such as Li et al. [102] propose a multimodal sentiment analysis method based on multi-level correlation mining and self-supervised multi-task learning, Peng et al. [103] propose a fine-grained modal label-based multi-stage network (FmlMSN), which utilize seven sentiment labels in unimodal, bimodal and trimodal information at different granularities from text, audio, image and the combinations of them. Researchers generally focus on the scale-level modal alignment and modal fusion before model’ decision. Sharafi et al. [96] design a new fusion method was proposed for multimodal emotion recognition utilizing different scales.
V-C Multimodal Alignment
Multimodal alignment involves synchronizing modal semantics before fusing multimodal data. A key challenge is handling missing modalities, which can occur due to issues like a camera being turned off, a user being silent, or device errors affecting both voice and text. Since the assumption of always having all modalities is often unrealistic, multimodal alignment must address these gaps. Additionally, it involves aligning objects across images, text, and audio through semantic alignment. Thus, we discuss multimodal alignment in terms of managing missing modalities and achieving semantic alignment. Fig. 3 illustrates the multimodal alignment.
V-C1 Alignment for Missing Modality
In real-world scenarios, data collection can sometimes result in the simultaneous loss of certain modalities due to unforeseen events. While multimodal affective computing typically assumes the availability of all modalities, this assumption often fails in practice, which can cause issues in modal fusion and alignment models when some modalities are missing. We classify existing methods for handling missing modalities into four groups.
The first group features the data augmentation approach, which randomly ablates the inputs to mimic missing modality cases. Parthasarathy et al. [110] propose a strategy to randomly ablate visual inputs during training at the clip or frame level to mimic real world scenarios. Wang et al. [111] deal with the utterance-level modalities missing problem by training emotion recognition model with iterative data augmentation by learned common representation. The second group is based on generative methods to directly predict the missing modalities given the available modalities [133]. For example, Zhao et al. [109] propose a missing modality imagination network (MMIN), which can predict the representation of any missing modality given available modalities under different missing modality conditions, so as to to deal with the uncertain missing modality problem. Zeng et al. [112] propose an ensemble-based missing modality reconstruction (EMMR) network to detect and recover semantic features of the key missing modality. Yuan et al. [113] propose a transformer-based feature reconstruction network (TFR-Net), which improves the robustness of models for the random missing in non-aligned modality sequences. Luo et al. [114] propose the multimodal reconstruction and align net (MRAN) to tackle the missing modality problem, especially to relieve the decline caused by the text modality’s absence.
The third group aims to learn the joint multimodal representations that can contain related information from these modalities [134]. For example, Ma et al. [135] propose a unified deep learning framework to efficiently handle missing labels and missing modalities for audio-visual emotion recognition through correlation analysis. Zeng et al. [116] propose a tag-assisted Transformer encoder (TATE) network to handle the problem of missing uncertain modalities, which designs a tag encoding module to cover both the single modality and multiple modalities missing cases, so as to guide the network’s attention to those missing modalities. Zuo et al. [117] propose to use invariant features for a missing modality imagination network (IF-MMIN), which includes an invariant feature learning strategy and an invariant feature based imagination module (IF-IM). Through the two strategies, IF-MMIN can alleviate the modality gap during the missing modalities prediction, thus improving the robustness of multimodal joint representation. Zhou et al. [119] propose a novel brain tumor segmentation network in the case of missing one or more modalities. The proposed network consists of three sub-networks: a feature-enhanced generator, a correlation constraint block and a segmentation network. The last group is translation-base methods. Tang et al. [101] propose the coupled-translation fusion network (CTFN) to model bi-direction interplay via couple learning, ensuring the robustness in respect to missing modalities. Liu et al. [118] propose a modality translation-based MSA model (MTMSA), which is robust to uncertain missing modalities. In summary, the works about alignment for miss modality focus on miss modality reconstruction and learning based on the available modal information.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.8em,font=, where level=2text width=6.3em,font=, where level=3text width=5.5em,font=, [ Multimodal Affective Computing, ver [ Multitask
Learning (§VI-A) [ MSA (§VI-A1) [ Self-MM [136], ARGF [137], MultiSE [138], VCAN [139], DTN [140], MMMIE [90], MMIM [88], MISA [91], , leaf, text width=34.1em ] ] [ MERC (§VI-A2) [ FacialMMT [25], MMMIE [90], AuxEmo [141], TDFNet [142], MALN [143], LGCCT [144], MultiEMO [84],
RLEMO [145], ,leaf, text width=34.1em ] ] [ MABSA (§VI-A3) [ CMMT [146], AbCoRD [147], JML [148], MPT [37], MMRBN [85], ,leaf, text width=34.1em ] ] [ MMER (§VI-A4) [ AMP [95], MEGLN-LDA [149], MultiSE [138], AMP [95]. ,leaf, text width=34.1em ] ] ] [ Pre-trained
Model (§VI-B) [ MSA (§VI-B1) [ MAG-XLNet [22], UniMSE [38], AOBERT [150], SKESL [151], TEASAL [152], TO-BERT [153], SPT [154],
ALMT [155] , leaf, text width=34.1em ] ] [ MERC (§VI-B2) [ FacialMMT [25], QAP [20], UniMSE [38], GraphSmile [156], ,leaf, text width=34.1em ] ] [ MABSA (§VI-B3) [ MIMN [25], GMP [18], ERUP [157], VLP-MABSA [158], DR-BERT [159], DTCA [160], MSRA [161],
AOF-ABSA [162], AD-GCFN [163], MOCOLNet [164],, ,leaf, text width=34.1em ] ] [ MMER (§VI-B4) [ TAILOR [94], ,leaf, text width=34.1em ] ] ] [ Enhanced
Knowledge
VI-C) [ MSA (§VI-C1) [ TETFN [165], ITP [19], SKEAFN [166], SAWFN [167], MTAG [107], ,leaf, text width=34.1em ] ] [ MERC (§VI-C2) [ ConSK-GCN [168], DMD [169], MRST [170], SF [171], TGMFN [172], RLEMO [145], DEAN [173], ,leaf, text width=34.1em ] ] [ MABSA (§VI-C3) [ KNIT [174], FITE [175], CoolNet [176], HIMT [177], ,leaf, text width=34.1em ] ] [ MMER (§VI-C4) [ UniVA-RoBERTa [178], CARAT [179], M3TR [180], MAGDRA [86], HHMPN [181], ,leaf, text width=34.1em ] ] ] [ Contextual
Information
VI-D) [ MSA (§VI-D1) [ MuLT [77], CIA [182], CAT-LSTM [183], CAMFNet [184], MTAG [107], CTNet [185], ScaleVLAD [104],
MMML [186], GFML [186], CHFusion [108], , leaf, text width=34.1em ] ] [ MERC (§VI-D2) [ CMCF-SRNet [83], MMGCN [187], MM-DFN [188], SAMGN [189], M3Net [190], M3GAT [187], RL-EMO [145],
SCMFN [191], EmoCaps [192], GA2MIF [193], MALN [143], COGMEN [49], ,leaf, text width=34.1em ] ] [ MABSA (§VI-D3) [ DTCA [160], MCPR [194], Elbphilharmonie [195], M2DF [196], AoM [197], FGSN [198], MIMN [16], ,leaf, text width=34.1em ] ] [ MMER (§VI-D4) [ MMS2S [199], MESGN [200], MDI [201], ,leaf, text width=34.1em ] ] ] ]
Figure 4: Taxonomy of multimodal affective computing works from aspects multitask learning, pre-trained model, enhanced knowledge and contextual information.
V-C2 Alignment for Cross-modal Semantics
Semantic alignment aims to find the connection between multiple modalities in one sample, which refers to searching one modal information through another modal information and vice versa. In the filed of MSA, Tsai et al. [77] leverage cross-modality and multi-scale modal alignment to implement the modal consistency in the semantic aspects, respectively. ScaleVLAD [202] proposes a fusion model to gather multi-Scale representation from text, video, and audio with shared vectors of locally aggregated descriptors to improve unaligned multimodal sentiment analysis. Yang et al. [107] convert unaligned multimodal sequence data into a graph with heterogeneous nodes and edges that captures the rich interactions across modalities and through time. Lee et al. [203] segment the audio and the underlying text signals into equal number of steps in an aligned way so that the same time steps of the sequential signals cover the same time span in the signals. Zong et al. [204] exploit multiple bi-direction translations, leading to double multimodal fusing embeddings compared with traditional translation methods. Wang et al. [205] propose a multimodal encoding–decoding translation network with a transformer and adopted a joint encoding–decoding method with text as the primary information and sound and image as the secondary information. Zhang et al. [123] propose a novel multi-level alignment to bridge the gap between acoustic and lexical modalities, which can effectively contrast both the instance-level and prototype-level relationships, separating the multimodal features in the latent space. Yu et al. [206] propose an unsupervised approach which minimizes the Wasserstein distance between both modalities, forcing both encoders to produce more appropriate representations for the final extraction to align text and image. Lai et al. [122] propose a deep modal shared information learning module based on the covariance matrix to capture the shared information between modalities. Additionally, we use a label generation module based on a self- supervised learning strategy to capture the private information of the modalities. Our module is plug-and-play in multimodal tasks, and by changing the parameterization, it can adjust the information exchange relationship between the modes and learn the private or shared information between the specified modes. We also employ a multi-task learning strategy to help the model focus its attention on the modal differentiation training data. For model robustness, Robust-MSA [121] present an interactive platform that visualizes the impact of modality noise to help researchers improve model capacity.
VI Models Across Multimodal Affective Computing
Figure 5: Illustration of multitask learning in multimodal affective computing tasks.
In the community of multimodal affective computing, the works appear to significant consistency in term of development technical route. For clarity, we group the these works based on multitask learning, pre-trained model, enhanced knowledge, contextual information. Meanwhile, we briefly summarized the advancements of MSA, MERC, MABSA and MMER tasks through the above four aspects. Fig. 4 summarizes the typical works of multimodal affective computing from these aspects and Table II shows the taxonomy of multimodal affective computing.
VI-A Multitask Learning
Multitask learning trains a model on multiple related tasks simultaneously, using shared information to enhance performance. The loss function combines losses from all tasks, with model parameters updated via gradient descent. In multimodal affective computing, multitask learning helps distinguish between modal-invariant and modal-specific features and integrates emotion-related sub-tasks into a unified framework. Fig. 5 shows the learning paradigm of multitask learning in multimodal affective learning task.
VI-A1 Multimodal Sentiment Analysis
In filed of multimodal sentiment analysis, Self-MM [136] generates a pseudo-label [207, 208, 209] for single modality and then jointly train unimodal and multimodal representations based on the generated and original labels. Furthermore, a translation framework ARGF between modalities, i.e., translating from one modality to another is used as an auxiliary task to regualize the multimodal representation learning [137]. Akhtar et al. [138] leverage the interdependence of the tasks sentiment and emotion to improve the model performance on two tasks. Chen et al. [139] propose a video-based cross-modal auxiliary network (VCAN), which is comprised of an audio features map module and a cross-modal selection module to make use of auxiliary information. Zheng et al. [140] propose a disentanglement translation network (DTN) with slack reconstruction to capture desirable information properties, obtain a unified feature distribution and reduce redundancy. Zheng et al. [90] combine mutual information maximization (MMMIE) between modal pairs with mutual information minimization between input data and corresponding features, jointly extract modal-invariant and task-related information in a single architecture.
VI-A2 Multimodal Emotion Recognition in Conversation
In community of multimodal emotion recognition, Zheng et al. [25] propose a two-stage framework named Facial expression-aware multimodal multi-task learning (FacialMMT), which jointly trains multimodal face recognition, unsupervised face clustering, and face matching in a unified architecture, so as to leverages the frame-level facial emotion distributions to help improve utterance-level emotion recognition based on multi-task learning. Zhang et al. [210] design two kinds of multitask learning (MTL) decoders, i.e., single-level and multi-level decoders, to explore their potential. More specifically, the core of a single-level decoder is a masked outer-modal self-attention mechanism. Sun et al. [141] design two auxiliary tasks to alleviate the insufficient fusion between modalities and guide the network to capture and align emotion-related features. Zhao et al. [142] propose a transformer-based deep-scale fusion network (TDFNet) for multimodal emotion recognition, solving the aforementioned problems. The multimodal embedding (ME) module in TDFNet uses pre-trained models to alleviate the data scarcity problem by providing a prior knowledge of multimodal information to the model with the help of a large amount of unlabeled data. Ren et al. [143] propose a novel multimodal adversarial learning network (MALN), which first mines the speaker’s characteristics from context sequences and then incorporate them with the unimodal features. Liu et al. [144] propose LGCCT, a light gated and crossed complementation transformer for multimodal speech emotion recognition.
VI-A3 Multimodal Aspect-based Sentiment Analysis
Yang et al. [146] propose a multi-task learning framework named cross-modal multitask Transformer (CMMT), which incorporates two auxiliary tasks to learn the aspect/sentiment-aware intra-modal representations and introduces a text-guided cross-modal interaction module to dynamically control the contributions of the visual information to the representation of each word in the inter-modal interaction. Jain et al. [147] propose a hierarchical multimodal generative approach (AbCoRD) for aspect-based complaint and rationale detection that reframes the multitasking problem as a multimodal text-to-text generation task. Ju et al. [148] is the first to jointly perform multimodal ATE (MATE) and multimodal ASC (MASC), and propose a joint framework JML with auxiliary cross-modal relation detection for multimodal aspect-level sentiment analysis (MALSA) to control the proper exploitation of visual information. Zou et al. [37] design a multimodal prompt Transformer (MPT) to perform cross-modal information fusion. Meanwhile, this work used the hybrid contrastive learning (HCL) strategy to optimize the model’s ability to handle labels with few samples. Chen et al. [85] design that audio module should be more expressive than the text module, and the single-modality emotional representation should be dynamically fused into the multimodal emotion representation, and proposes corresponding rule-based multimodal multi-task network (MMRBN) to restrict the representation learning.
VI-A4 Multimodal Multi-label Emotion Recognition
For multimodal multi-label emotion recognition, Ge et al. [95] design adversarial temporal masking strategy and adversarial parameter perturbation strategy to jointly enhance the encoding of other modalities and generalization of the model respectively. MER-MULTI [149] is label distribution adaptation which adapts the label distribution between the training set and testing set to remove training samples that do not match the features of the testing set. Akhtar et al. [211] present a deep multi-task learning framework that jointly performs sentiment and emotion analysis both, which leverage the inter-dependence of two related tasks (i.e. sentiment and emotion) in improving each others performance using an effective multimodal framework.
VI-B Pre-trained Model
In recent years, large language model (LLM) [59, 212] and multimodal pre-trained model [213, 27, 22, 214] has achieved significant progress [26, 212, 215]. Compared with non-pretrained model, pre-trained model contains massive transferred knowledge [32, 28], which can be introduced into multimodal representation learning to probe the richer information. Fig. 6 shows the use of pre-trained model in multimodal affective learning task.
Figure 6: An illustration of pre-trained model in multimodal affective computing tasks. PLM denotes pre-trained language model.
VI-B1 Multimodal Sentiment Analysis
In filed of multimodal sentiment analysis, Rahman et al. [22] propose an attachment to pre-trained model BERT and XLNet called multimodal adaptation gate (MAG), which allows BERT and XL-Net to accept multimodal nonverbal data by generating a shift that is conditioned on the visual and acoustic modalities to internal representation of BERT and XLNet. UniMSE [38] is a unified sentiment-sharing framework based on T5 model [60], which injects the non-verbal signals into pre-trained Transformer-based model for probing the knowledge stored in LLM. AOBERT [150] introduces a single-stream transformer structure, which integrates all modalities into one BERT model. Qian et al. [151] embed sentiment information at the word level into pre-trained multimodal representation to facilitate further learning on limited labeled data. TEASAL [152] is a Transformer-Based speech-prefixed language model, which exploits a conventional pre-trained language model as a cross-modal Transformer encoder. Yu et al. [153] study target-oriented multimodal sentiment classification (TMSC) and propose a multimodal BERT architecture for multimodal sentiment analysis task. Cheng et al. [154] set layer-wise parameter sharing and factorized co-attention that share parameters between cross attention blocks, so as to allow multimodal signal to interact within every layer. ALMT [155] incorporates an adaptive hyper-modality learning (AHL) module to learn an irrelevance/conflict-suppressing representation from visual and audio features under the guidance of language features at different scales.
VI-B2 Multimodal Emotion Recognition in Conversation
In the domain of multimodal emotion recognition in conversation, FacialMMT [25] is a two-stage framework, which takes RoBERTa [216] and Swin Transformer as the backbone for representation learning. Qiu et al. [217] adopt VATT [31] to encode vision, text and audio respectively, and makes an alignment between the learned modal representation. QAP [20] is a quantum-inspired adaptive-priority-learning model, which employs ALBERT as the text encoder and introduces quantum theory (QT) to learn modal priority adaptively. UniMSE [38] proposes a multimodal fusion method based on pre-trained model T5, aiming to fuse the modal information with pre-trained knowledge. GraphSmile [156] adopts Roberta [216] to track intricate emotional cues in multimodal dialogues, alternately assimilating inter-modal and intra-modal emotional dependencies layer by layer, adequately capturing cross-modal cues while effectively circumventing fusion conflicts.
TABLE II: Taxonomy of multimodal affective computing task from tasks of MSA, MERC, MABSA and MMER.
Task Model Model architecture Miss Modality Datasets
MSA TFN [76] LSTM MOSI,MOSEI
MuLT [77] cross-attention MOSI,MOSEI
MISA [91] Transformer MOSI,MOSEI,UR_FUNNY
Self-MM [136] BiLSTM MOSI,MOSEI,SIMS
MTAG [107] Graph-based neural network MOSI, IEMOCAP
ScaleVLAD [104] VLAD,CNN MOSI,MOSEI,IEMOCAP
MMIM [88] BERT, CPC MOSI,MOSEI
UniMSE [38] Adapter, T5, contrastive learning MOSI,MOSEI,IEMOCAP,MELD
CHFusion [108] CNN MOSI,IEMOCAP
MMHA [218] Multi-head Attention MOUD,MOSI
QMR [219] Quantum Language Mode Getty Images,Flickr
RMFN [100] LSTHM MOSI
HMM-BLSTM [220] HMM,BiLSTM IEMOCAP
HALCB [221] Cognitive brain limbic system MOSI,YouTube,MOSEI
CSFC [222] CNN,fuzzy logic Alh,Mos,Sag,MOUD
SFNN [223] CNN,attention vista-net
GFML [186] Multi-head attention MOSEI,MOSI
MMML [186] cross-attention MOSEI,MOSI,CH-SIMS
TATE [116] Transformer IEMOCAP,MOSI
MERC FacialMMT-RoBERTa [224] RoBERTa MELD, Aff-Wild2
MultiEMO [84] Multi-head attention IEMOCAP, MELD
MMGCN [187] Graph convolutional network IEMOCAP, MELD
MM-DFN [188] GRU,Graph conventional network IEMOCAP,MELD
EmoCaps [192] Multi-head attention IEMOCAP,MELD
GA2MIF [193] Graph attention networks,Multi-head attention IEMOCAP,MELD
MALN [143] LSTM, cross-attention, Transformer-based model IEMOCAP, MELD
M2R2 [111] GRU IEMOCAP, MELD
TDF-Net [142] CNN,GRU,Transformer IEMOCAP
SDT [225] CNN,Transformer IEMOCAP,MELD
HCT-DMG [226] Cross-modal Transformer CMU-MOSI, MOSEI,IEMOCAP
COGMEN [49] Graph neural network, Transformer CMU-MOSI, IEMOCAP
Qiu et al. [217] VATT(Video-Audio-Text Transformer) CMU-MOSEI
QAP [20] VGG,AlBERT IEMOCAP,CMU-MOSEI
MPT-HCL [37] Transformer IEMOCAP,MELD
CMCF-SRNet [83] Transformer,RGCN IEMOCAP,MELD
SAMGN [189] Graph neural network IEMOCAP,MELD
M3Net [190] Graph neural network IEMOCAP,MELD
M3GAT [227] Graph attention network MELD,MEISD,MESD
IF-MMIN [117] Imagination Module IEMOCAP
AMuSE [87] Attention-based model IEMOCAP,MELD
DEAN [173] Transformer MOSI,MOSEI,IEMOCAP
MMIN [109] Transformer IEMOCAP,MSP-IMPROV
RLEMO [145] Graph Convolutional Network,GRU IEMOCAP,MELD
Yao et al. [191] Graph convolutional network IEMOCAP,MELD
BCFN [228] CNN CHERMA,CH-SIMS
MM-RBN [85] Transformer IEMOCAP
MABSA AoM [197] BART,CNN Twitter2015, Twitter2017
JML [148] BERT, ResNet TRC, Twitter2015, Twitter2017
VLP-MABSA [158] Vision-Language pre-trained model Twitter2015, Twitter2017
CMMT [146] cross-attention,self-attention Twitter-2015,Twitter-2017,Political-Twitter
DTCA [160] RoBERTa, ViT Twitter2015, Twitter2017
M2DF [229] pre-trained model CLIP Twitter2015, Twitter2017
MIMN [16] Memory Network ZOL
KNIT [174] Transformer, Twitter2015, Twitter2017
MMER MMS2S [199] cross-attention, multi-head attention MOSEI
MESGN [200] cross-modal transformer MOSEI
TAILOR [230] Transformer, cross-attention MOSEI
HHMPN [231] MPNN, multi-head attention MOSEI
AMP [95] Transformer-based encoder-decoder, mask learning CMU-MOSEI,NEMu
M3TR [232] CNN, Transformer, cross-attention MS-COCO,VOC 2007
UniVA [178] VAD model, contrastive learning MOSEI, $M^{3}ED$
VI-B3 Multimodal Aspect-based Sentiment Analysis
In the study of multimodal aspect-based sentiment analysis, Xu et al. [50] are the first to put forward the task, multimodal aspect-based sentiment analysis, and propose a novel multi-interactive memory network (MIMN), which includes two interactive memory networks to supervise the textual and visual information with the given aspect, and learns not only the interactive influences between cross-modality data but also the self influences in single-modality data. Yang et al. [18] propose a novel generative multimodal prompt (GMP) model for MABSA, which includes the multimodal encoder module and the N-Stream decoders module and perform three MABSA-related tasks with quite a small number of labeled multimodal samples. Liu et al. [157] propose an entity-related unsupervised pre-training with visual prompts for MABSA. Instead of using sentiment-related supervised pre-training, two entity-related unsupervised pre-training tasks are applied and compared, which are targeted at locating the entities in text with the support of visual prompts. Ling et al. [158] propose a task-specific Vision-Language Pre-training framework for MABSA (VLP-MABSA), which is a unified multimodal encoder-decoder architecture for all the pre-training and downstream tasks. Zhang et al. [159] construct a dynamic re-weighting BERT (DR-BERT) based BERT and designed to learn dynamic aspect-oriented semantics for ABSA. Jin et al. [161] propose a multi-aspect semantic relevance model that takes into account the match between search queries and the title, attribute and image information of items simultaneously. Wang et al. [162] propose an end-to-end MABSA framework with image conversion and noise filtration, which bridges the representation gap in different modalities by translating images into the input space of a pre-trained language model (PLM). Wang et al. [163] propose an adaptive dual graph convolution fusion network (AD-GCFN) for aspect-based sentiment analysis. This model uses two graph convolution networks: one for the semantic layer to learn semantic correlations by an attention mechanism, and the other for the syntactic layer to learn syntactic structure by dependency parsing. Mu et al. [164] propose a novel momentum contrastive learning network (MOCOLNet) which is an unified multimodal encoder-decoder framework for multimodal aspect-level sentiment analysis task. An end-to-end training manner is proposed to optimize MOCOLNet parameters, which alleviates the problem of scarcity of labelled pre-training data. To fully explore multimodal contents especially the location information, Zhang et al. [233] design a multimodal interactive network to model the textual and visual information of each aspect. Two memory network-based modules are used to capture intra-modality features, including text to aspect alignment and image to aspect alignment. Yang et al. [234] propose multi-grained fusion network with self-distillation (MGFN-SD) to analyze aspect-based sentiment polarity, which can effectively integrate multi-grained representation learning with self-distillation to obtain more representative multimodal features.
VI-B4 Multimodal Multi-label Emotion Recognition
A few works on multimodal multi-label emotion recognition leverage pre-trained model to improve model performance. To our best known, TAILOR [94] is a novel framework of versatile multimodal learning for multi-labeL emotion recognition, which adversarially depicts commonality and diversity among multiple modalities. TAILOR adversarially extracts private and common modality representations. Then a BERT-like transformer encoder is devised to gradually fuse these representations in a granularity descent way.
VI-C Enhanced Knowledge
Figure 7: Illustration of enhanced knowledge in multimodal affective computing tasks.
External knowledge in machine learning and AI refers to information from outside the training dataset, including knowledge bases, text corpora, knowledge graphs, pre-trained models, and expert insights. Integrating this knowledge can improve performance, generalization, interpretability, and robustness to noisy or limited data [235, 236]. Fig. 7 shows the common way of incorporating external knowledge into multimodal affective learning task.
VI-C1 Multimodal Sentiment Analysis
In the area of study focused on multimodal sentiment analysis, Rahmani et al. [19] construct an adaptive tree by hierarchically dividing users and utilizes an attention-based fusion to transfer cognitive-oriented knowledge within the tree. TETFN [165] is a novel method named text enhanced Transformer fusion network, which learns text-oriented pairwise cross-modal mappings for obtaining effective unified multimodal representations. Zhu et al. [166] propose the sentiment knowledge enhanced attention fusion network (SKEAFN), a novel end-to-end fusion network that enhances multimodal fusion by incorporating additional sentiment knowledge representations from an external knowledge base. Chen et al. [167] try to incorporate sentimental words knowledge into the fusion network to guide the learning of joint representation of multimodal features.
VI-C2 Multimodal Emotion Recognition in Conversation
In the discipline of research related to multimodal emotion recognition in conversation, Fu et al. [168] integrate context modeling, knowledge enrichment, and multimodal (text and audio) learning into a GCN-based architecture. Li et al. [169] propose a decoupled multimodal distillation (DMD) approach that facilitates flexible and adaptive crossmodal knowledge distillation, aiming to enhance the discriminative features of each modality. Sun et al. [170] investigate a multimodal fusion transformer network based on rough set theory, which facilitates the interaction of multimodal information and feature guidance through rough set cross-attention. Wang et al. [171] devise a novel label-guided attentive fusion module to fuse the label-aware text and speech representations, which learns the label-enhanced text/speech representations for each utterance via label-token and label-frame interactions.
VI-C3 Multimodal Aspect-based Sentiment Analysis
In the field of research on multimodal aspect-based sentiment analysis, Xu et al. [174] introduce external knowledge, including textual syntax and cross-modal relevancy knowledge to Transformer layer, which cut off the irrelevant connections among textual or cross-modal modalities by using a knowledge-induced matrix. Yang et al. [175] distill visual emotional cues and align them with the textual content to selectively match and fuse with the target aspect in textual modality. CoolNet [176] is a cross-modal fine-grained alignment and fusion network, aims to boost the performance of visual-language models in seamlessly integrating vision and language information, which transforms an image into a textual caption and a graph structure, then dynamically aligns the semantic and syntactic information from both the input sentence and the generated caption, as well as models the object-level visual features. To strengthen the semantic meanings of image representations, Yu et al. [177] propose to detect a set of salient objects in each image based on a pre-trained Faster R-CNN model, and represent each object by concatenating its hidden representation and associated semantic concepts, followed by an Aspect-Guided Attention layer to learn the relevance of each semantic concept with the guidance of given aspects.
VI-C4 Multimodal Multi-label Emotion Recognition
In the area of study focused on multimodal multi-label emotion recognition, Zheng et al. [178] propose to represent each emotion category with the valence-arousal (VA) space to capture the correlation between emotion categories and design a unimodal VA-driven contrastive learning algorithm. CARAT [179] presents contrastive feature reconstruction and aggregation for the MMER task. Specifically, CARAT devises a reconstruction-based fusion mechanism to better model fine-grained modality-to-label dependencies by contrastively learning modal-separated and label-specific features. Zhao et al. [180] propose a novel multimodal multi-label TRansformer (M3TR) learning framework, which embeds the high-level semantics, visual structures and label-wise co-occurrences of multiple modalities into one unified encoding. Li et al. [86] propose a novel multimodal attention graph network with dynamic routing-by-agreement (MAGDRA) for MMER. MAGDRA is able to efficiently fuse graph data with various node and edge types as well as properly learn the cross-modal and temporal interactions between multimodal data without pre-aligning. Zhang et al. [181] propose heterogeneous hierarchical message passing network (HHMPN), which can simultaneously model the feature-to-label, label-to-label and modality-to-label dependencies via graph message passing.
VI-D Contextual Information
Figure 8: Illustration of context information in multimodal affective computing tasks.
Context refers to the surrounding words, sentences, or paragraphs that give meaning to a particular word or phrase. Understanding context is crucial for tasks like dialogue systems or sentiment analysis. In a conversation, context includes the history of previous utterances, while for news, it refers to the overall description provided by the entire document. Overall, contextual information helps machines make more accurate predictions. Fig. 8 shows the significance of context information to multimodal affective learning task.
VI-D1 Multimodal Sentiment Analysis
In the community of multimodal sentiment analysis, Chauhan et al. [182] employ a context-aware attention module to learn intra-modality interaction among particating modalities through encoder-decoder structure. Multimodal context integrate the unimodal context, Poria et al. [183] propose a recurrent model with multi-level multiple attentions to capture contextual information among utterances, and design a recurrent model to capture contextual information among utterances and introduced attention-based networks for improving both context learning and dynamic feature fusion. Huang et al. [184] propose a novel context-based adaptive multimodal fusion network (CAMFNet) for consecutive frame-level sentiment prediction. Li et al. [185] propose a spatial context extraction block to explore the spatial context by calculating the relationships between feature maps and the higher-level semantic representation in images.
VI-D2 Multimodal Emotion Recognition in Conversation
In the realm of research concerning multimodal emotion recognition in conversation, Hu et al. [187] make use of multimodal dependencies effectively, and leverages speaker information to model inter-speaker and intra-speaker dependency. Zhang et al. [83] propose a cross-modality context fusion and semantic refinement network (CMCF-SRNet) to solve the limitation of insufficient semantic relationship information between utterances. Zhang et al. [189] construct multiple modality-specific graphs to model the heterogeneity of the multimodal context. Chen et al. [190] propose a GNN-based model that explores multivariate relationships and captures the varying importance of emotion discrepancy and commonality by valuing multi-frequency signals. Zhang et al. [227] propose a multimodal, multi-task interactive graph attention network, termed M3GAT, to simultaneously solve conversational context dependency, multimodal interaction, and multi-task correlation in a unified framework. RL-EMO [145] is a novel reinforcement Learning framework for the multimodal emotion recognition task, which combines reinforcement learning (RL) module to model context at both the semantic and emotional levels respectively. Yao et al. [191] propose a speaker-centric multimodal fusion network for emotion recognition in a conversation, to model intra-modal feature fusion and speaker-centric cross-modal feature fusion.
VI-D3 Multimodal Aspect-based Sentiment Analysis
In the stduy of multimodal aspect-based sentiment analysis, Yu et al. [160] propose an unsupervised approach which minimizes the Wasserstein distance between both modalities, forcing both encoders to produce more appropriate representations for the final extraction. Xu et al. [194] design and construct a multimodal Chinese product review dataset (MCPR) to support the research of MABSA. Anschutz et al. [195] report the results of an empirical study on how semantic computing can provide insights into user-generated content for domain experts. In addition, this work discussed different image-based aspect retrieval and aspect-based sentiment analysis approaches to handle and structure large datasets. Zhao et al. [196] borrow the idea of Curriculum Learning and propose a multi-grained multi-curriculum denoising Framework (M2DF) to adjust the order of training data, so as to obtain more contextual information. Zhou et al. [197] propose an aspect-oriented method (AoM) to detect aspect-relevant semantic and sentiment information. Specifically, an aspect-aware attention module is designed to simultaneously select textual tokens and image blocks that are semantically related to the aspects. Zhao et al. [198] propose a fusion with GCN and SE ResNeXt Network (FGSN), which constructs a graph convolution network on the dependency tree of sentences to obtain the context representation and aspects words representation by using syntactic information and word dependency.
VI-D4 Multimodal Multi-label Emotion Recognition
MMS2S [199] is a multimodal sequence-to-set approach to effectively model label dependence and modality dependence. MESGN [200] firstly proposes this task, which simultaneously models the modality-to-label and label-to-label dependencies. Many works consider the dependencies of multi-label based on the characteristics of co-occurrence labels. Zhao et al. [201] propose a general multimodal dialogue-aware interaction framework, named by MDI, to model the impacts of dialogue context on emotion recognition.
TABLE III: List of multimodal affective computing datasets. T,A,V denote the text, audio and vision modalities, respectively. Emotion denotes the sample in dataset is labeled with emotion category and sentiment denotes the sample in dateset is labeled with sentiment polarity.
Task Dataset Modalities Source Emotion Sentiment Language Datasize
MSA CMU-MOSI [237] T,A,V Video Blog, YouTube English 2,199
CMU-MOSEI [238] T,A,V YouTube English 22,856
CH-SIMS [239] T,A,V Movies,TVs Chinese 2,281
CMU-MOSEAS [240] T,A,V YouTube Spanish,Portuguese,German,French 4,000
ICT-MMMO [241] T,A,V reviews - - - -
YouTube [46] T,A,V YouTube - - English 300
CH-SIMS v2.0[242] T,A,V TV series,Shows,Movies Chinese 14,402
MERC MELD [243] T,A,V Friends TV English 13,707
IEMOCAP [244] T,A,V Act English 7,532
HED [245] T,V Movies, TVs English 17,441
RML [246] A,V Video English, Mandarin, Urdu, Punjabi, Persian, and Italian -
BAUM-1 [247] A,V Data collection Turkish 1,184
MAHNOB-HCI[248] V, EEG Data collection - -
Deap [249] EEG Act, Data collection physiological signal -
MuSe-CaR [250] T,A,V Car reviews,YouTube English
CHEAVD [251] A,V Movies,TVs - - Mandarin 7,030
MSP-IMPROV [252] T,A,V Act English 8,438
MEISD [253] T,A,V TVs English -
MESD [254] T,A,V TVs English 9,190
Ulm-TSST [255] A,V,EEG job interviews English -
CHERMA [256] T,A,V TV series,Shows,Movies Chinese 28,717
AMIGOS [257] EEG, ECG, GSR Data collection - -
MABSA Twitter2015 [258] T,V Twitter English -
Twitter2017 [258] T,V Twitter English -
MCPR [259] T,V Product reviews Chinese 15,000
Multi-ZOL [50] T,V Product reviews Chinese 5,288
MACSA [260] T,V Hotel service reviews Chinese 21,108
MASAD [261] T,V Visual sentiment ontology datasets English 38,532
PanoSent [262] T,A,V Social media English,Chinese, Spanish 10,000
MMER CMU-MOSEI [238] T,A,V YouTube English 22,856
$M^{3}ED$ [201] T,A,V 56 TVs Mandarin 24,449
VII Datasets of Multimodal Affective Computing
In this section, we introduce the benchmark datasets of MSA, MERC, MABSA, and MMER tasks. To facilitate easy navigation and reference, the details of datasets are shown in Table III with a comprehensive overview of the studies that we cover.
VII-A Multimodal Sentiment Analysis
  • MOSI [237] contains 2,199 utterance video segments, and each segment is manually annotated with a sentiment score ranging from -3 to +3 to indicate the sentiment polarity and relative sentiment strength of the segment.
  • MOSEI [238] is an upgraded version of MOSI, annotated with both sentiment and emotion. MOSEI contains 22,856 movie review clips from YouTube. Each sample in MOSEI includes sentiment annotations ranging from -3 to +3 and multi-label emotion annotations.
  • CH-SIMS [239] is a Chinese single- and multimodal sentiment analysis dataset, which contains 2,281 refined video segments in the wild with both multimodal and independent unimodal annotations.
  • CH-SIMS v2.0 [242] is an extended version of CH-SIMS that includes more data instances, spanning text, audio and visual modalities. Each modality of sample is annotated with sentiment polarity, and then sample is annotated with a concluded sentiment.
  • CMU-MOSEAS [240] is the first large-scale multimodal language dataset for Spanish, Portuguese, German and French, and it is collect from YouTube and its samples are 4,000 in total.
  • ICT-MMMO [241] is collected from online social review videos that encompass a strong diversity in how people express opinions about movies and include a real-world variability in video recording quality [^4] http://multicomp.ict.usc.edu.
  • YouTube [46] collects 47 videos from the social media web site YouTube. Each video contains 3-11 utterances with most videos having 5-6 utterances in the extracted 30 seconds.
VII-B Multimodal Emotion Recognition in Conversation
  • MELD [243] contains 13,707 video clips of multi-party conversations, with labels following Ekman’s six universal emotions, including joy, sadness, fear, anger, surprise and disgust.
  • IEMOCAP [244] has 7,532 video clips of multi-party conversations, with labels following Ekman’s six universal emotions, including joy, sadness, fear, anger, surprise and disgust.
  • HED [245] contains happy, sad, disgust, angry and scared emotion-aligned face, body and text samples, which are much larger than existing datasets. Moreover, the emotion labels were correspondingly attached to those samples by strictly following a standard psychological paradigm.
  • RML [246] collects video samples from eight subjects, speaking six different languages. The six languages are English, Mandarin, Urdu, Punjabi, Persian, and Italian. This dataset contains 500 video samples, each delivered with one of the six particular emotions.
  • BAUM-1 [247] contains two sets: BAUM-1a and BAUM-1s databases. BAUM-1a database contains clips containing expressions of five basic emotions (happiness, sadness, anger, disgust, fear) along with expressions of boredom, confusion (unsure) and interest (curiosity). BAUM-1s database contains clips reflecting six basic emotions and also expressions of boredom, contempt, confusion, thinking, concentrating, bothered, and neutral.
  • MAHNOB-HCI [248] includes 527 facial video recordings of 27 participants engaged in various tasks and interactions, while their physiological signals such as 32-channel electroencephalogram (EEG), 3-channel electrocardiogram.
  • Deap [249] contains data from 32 participants, aged between 19 and 37 (50% female), who were recorded watching 40 one-minute music videos. Each participant was asked to evaluate each video by assigning values from 1 to 9 for arousal, valence, dominance, like/dislike, and familiarity.
  • MuSe-CaR [250] focuses on the tasks of emotion, emotion-target engagement, and trustworthiness recognition by means of comprehensively integrating the audio-visual and language modalities.
  • CHEAVD 2.0 [251] is selected from Chinese movies, soap operas and TV shows, which contains noise in the background to mimic real-world conditions.
  • MSP-IMPROV [252] is a multimodal emotional database comprised of spontaneous dyadic interactions, designed to study audiovisual perception of expressive behaviors.
  • MEISD [253] is a large-scale balanced multimodal multi-label emotion, intensity, and sentiment dialogue dataset (MEISD) collected from different TV series that has textual, audio, and visual features.
  • MESD [254] is the first multimodal and multi-task sentiment, emotion, and desire dataset, which contains 9,190 text-image pairs, with English text.
  • Ulm-TSST [255] is a multimodal dataset, where participants were recorded in a stressful situation emulating a job interview, following the TSST protocol.
  • CHERMA [256] provides uni-modal labels for each individual modality, and multi-modal labels for all modalities jointly observed. It is collected from various source, including 148 TV series, 7 variety shows, and 2 movies.
  • AMIGOS [257] is collected in two experimental settings. In the first setting, 40 participants viewed 16 short emotional videos. In the second setting, participants watched 4 longer videos, some individually and others in groups. During these sessions, participants’ physiological signals—Electroencephalogram (EEG), Electrocardiogram (ECG), and Galvanic Skin Response (GSR)—were recorded using wearable sensors.
VII-C Multimodal Aspect-based Sentiment Analysis
  • Twitter2015 and Twitter2017 are originally provided by the work [258] for multimodal named entity recognition and annotated with the sentiment polarity for each aspect by the work [14].
  • MCPR [259] has 2,719 text-image pairs and 610 distinct aspects in total, which collects 1.5k product reviews involving clothing and furniture departments, from the e-commercial platform JD.com. It is the first aspect-based multimodal Chinese product review dataset.
  • Multi-ZOL [50] consists of reviews of mobile phones collected from ZOL.com. It contains 5,288 sets of multimodal data points that cover various models of mobile phones from multiple brands. These data points are annotated with a sentiment intensity rating from 1 to 10 for six aspects.
  • MACSA [260] contains more than 21K text-image pairs, and provides fine-grained annotations for both textual and visual content and firstly uses the aspect category as the pivot to align the fine-grained elements between the two modalities.
  • MASAD [261] selects 38,532 samples from a partial VSO visual dataset [263] (approximately 120,000 samples) that can clearly express sentiments and categorized them into seven domains: food, goods, buildings, animal, human, plant, scenery, with a total of 57 predefined aspects.
  • PanoSent [262] is annotated both manually and automatically, featuring high quality, large scale (10,000 dialogues), multimodality (text, image, audio and video), multilingualism (English, Chinese and Spanish), multi-scenarios (over 100 domains), and covering both implicit&explicit sentiment elements.
VII-D Multimodal Multi-label Emotion Recognition
  • CMU-MOSEI [238] contains 22,856 movie review clips from Youtube videos. Each video intrinsically contains three modalities: text, audio, and visual, and each movie review clip is annotated with at least one emotion category of the set: angry, disgust, fear, happy, sad, surprise.
  • $\textbf{M}^{3}\textbf{ED}$ [201] is a multimodal emotional dialogue dataset in Chinese, which contains a total of 9,082 turns and 24,449 utterances, and each utterance is annotated with the seven emotion categories (happy, surprise, sad, disgust, anger, fear, and neutral).
VIII Evaluation Metrics
In this section, we report the mainstream evaluation metrics for each multimodal affective computing task.
Multimodal Sentiment Analysis
Previous works adopt mean absolute error (MAE), Pearson correlation (Corr), seven-class classification accuracy (ACC-7), binary classification accuracy (ACC-2) and F1 score computed for positive/negative and non-negative/negative classification as evaluation metrics.
Multimodal Emotion Recognition in a conversation
Accuracy (ACC) and weighted F1 (WF1) are used for evaluation. Additionally, the imbalance label distribution results in a phenomenon that the trained model performs better on some categories and perform poorly on others. In order to verify the impacts of data distribution on model performance, researchers also provide ACC and F1 on each emotion category to measure the model performance.
Multimodal Aspect-based Sentiment Analysis
With the previous methods, for multimodal aspect term extraction (MATE) and joint multimodal aspect sentiment analysis (JMASA) tasks, researchers use precision (P), recall (R) and micro-F1 (F1) as the evaluation metrics. For the multimodal aspect sentiment classification(MASC) task, accuracy (ACC) and macro-F1 are as evaluation metrics.
Multimodal Multi-label Emotion Recognition
According to the prior work, multi-label classification works mostly adopt accuracy (ACC), micro-F1, precision (P) and recall (R) as evaluation metrics.
IX Discuss
In this section, we briefly discuss the works of multimodal affective computing based on facial expression, acoustic signal, physiological signals, and emotion cause. Furthermore, we discuss the technical routes across multiple multimodal affective computing tasks to track their consistency and difference.
IX-A Other Multimodal Affective Computing
Multimodal Affective Computing Based on Facial Expression Recognition
Facial expression recognition has significantly evolved over the years, progressing from static to dynamic methods. Initially, static facial expression recognition (SFER) relied on single-frame images, utilizing traditional image processing techniques such as Local Binary Patterns (LBP) and Gabor filters to extract features for classification. The advent of deep learning brought Convolutional Neural Networks (CNNs), which markedly improved the accuracy of SFER [264, 265, 266, 267]. However, static methods were limited in capturing the temporal dynamics of facial expressions [268]. Some methods attempt to approach the problem from a local-global feature perspective, extracting more fine-grained visual representations and identifying key informative segments [269, 270, 271, 272, 273, 274]. These approaches enhance robustness against noisy frames, enabling uncertainty-aware inference. To further enhance accuracy, recent advancements in DFER focus on integrating multimodal data and employing parameter-efficient fine-tuning (PEFT) to adapt large pre-trained models for enhanced performance [275, 276, 277], while Liu et al. [278] introduces the concept of expression reenactment (i.e. normalization), harnessing generative AI to mitigate noise in in-the-wild datasets. Moreover, the burgeoning evidential deep learning (EDL) has shown considerable promise by enabling explicit uncertainty quantification through the distributional measurement in latent spaces for improved interpretability, with demonstrated efficacy in zero-shot learning [279], multi-view classification [280, 281, 282], video understanding [283, 284, 285] and multi-modal named entity recognition.
Multimodal Affective Computing Based on Acoustic Signal
The model based on single-sentence single-task is the most common model in speech emotion recognition. For example, Aldeneh et al. [286] use CNN to perform convolutions in the time direction of handcrafted temporal features (40-dimensional MFSC) to identify emotion-salient regions and used global max pooling to capture important temporal areas. Li et al. [287] apply two different convolution kernels on spectrograms to extract temporal and frequency domain features, concatenated them, and input them into a CNN for learning, followed by attention mechanism pooling for classification. Trigeorgis et al. [288] use CNN for end-to-end learning directly on speech signals, avoiding the problem of feature extraction not being robust for all speakers. Mirsamadi et al. [289] combine Bidirectional LSTM (Bi-LSTM) with a novel pooling strategy, utilizing attention mechanisms to enable the network to focus on emotionally prominent parts of sentences. Zhao et al. [290] consider the temporal and spatial characteristics of the spectrum in the attention mechanism to learn time-related features in spectrograms, and using CNN to learn frequency-related features in spectrograms. Luo et al. [291] propose a dual-channel speech emotion recognition model that uses CNN and RNN to learn from spectrograms on one hand, and separately learns HSFs features on the other, finally concatenating the obtained features for classification.
Multimodal Affective Computing Based on Physiological Signals
In medical measurements and health monitoring, EEG-based emotion recognition (EER) is one of the most promising directions within emotion recognition and has attracted substantial research attention [292, 293, 294]. Notably, the field of affective computing has seen nearly 1,000 publications related to EER since 2010 [295]. Numerous EEG-based multimodal emotion recognition (EMER) methods have been proposed [296, 297, 298, 299, 300], leveraging the complementarity and redundancy between EEG and other physiological signals in expressing emotions. For example, Vazquez et al. [301] address the problem of multimodal emotion recognition from multiple physiological signal, which demonstrates Transformer-based approach is suitable for emotion recognition based on physiological signal.
Multimodal Affective Computing Based on Emotion Cause
Apart from focusing on the emotions themselves, the capacity of machine for understanding the cause that triggers an emotion is essential for comprehending human behaviors, which makes emotion-cause pair extraction (ECPE) crucial. Over the years, text-based ECPE has made significant progress [302, 303]. Based on ECPE, Li et al. [304] propose multimodal emotion-cause pair extraction (MECPE), which aims to extract emotion-cause pairs with multimodal information. Initially, Li et al. [304] construct a joint training architecture, which contains the main task, i.e., multimodal emotion-cause pair extraction and two subtasks, i.e., multimodal emotion detection and cause detection. To solve MECPE, researchers borrowed the multitask learning framework to train the model using multiple training objectives of sub-tasks, aiming to enhance the knowledge sharing among them. For example, Li et al. [305] propose a novel model that captures holistic interaction and label constraint (HiLo) features for the MECPE task. HiLo enables cross-modality and cross-utterance feature interactions through various attention mechanisms, providing a strong foundation for accurate cause extraction.
IX-B Consistency among Multimodal Affective Computing
We categorize the multimodal affective computing tasks into several key areas: multimodal alignment and fusion, multi-task learning, pre-trained models, enhanced knowledge, and contextual information. To ensure clarity, we discuss the consistencies across these aspects.
Multimodal alignment and fusion
Among MSA, MERC, MABSA and MMER tasks, each is fundamentally a multimodal task that involves considering and combining at least two modalities to make decisions. This process includes extracting features from each modality and integrating them into a unified representation vector. In multimodal representation learning, modal alignment and fusion are two critical issues that must be addressed to advance the field of multimodal affective computing. For vision-dominated multimodal tasks such as image captioning [8, 306], the impact of vision is more significant than language. In contrast, multimodal affective computing tasks place a greater emphasis on language [38, 307].
Pre-trained model
Generally, pre-trained models are used to encode raw modal information into vectors. From this perspective, multimodal affective computing tasks adopt pre-trained models as the backbone and then fine-tune them for downstream tasks. For example, UniMSE [38] uses T5 as the backbone, while GMP [18] utilizes BART. These approaches aim to transfer the general knowledge embedded in pre-trained language models to the field of affective computing.
Enhanced knowledge
Commonsense knowledge encompasses facts and judgments about our natural world. In the field of affective computing, this knowledge is crucial for enabling machines to understand human emotions and their underlying causes. Researchers enhance affective computing by integrating external knowledge sources such as sentiment lexicons [308], English knowledge bases [309, 310, 311, 312, 313], and Chinese knowledge bases [314] as the external knowledge to enhance affective computing.
Contextual information
Affective computing tasks require an understanding of contextual information. In MERC, contextual information encompasses the entire conversation, including both previous and subsequent utterances relative to the current utterance. For MABSA, contextual information refers to the full sentence containing customer opinions. Researchers integrate contextual information using hierarchical approaches [315, 316], self-attention mechanisms [58], and graph-based dependency modeling [317, 318]. Additionally, affective computing tasks can enhance understanding by incorporating non-verbal cues such as facial expressions and vocal tone, alongside textual information.
IX-C Difference among Multimodal Affective Computing
We examine the differences among multimodal affective computing tasks by considering the type of downstream tasks, sentiment granularity, and application contexts to identify the unique characteristics of each task.
For downstream tasks, MSA predicts sentiment strength as a regression task. MERC is a multi-class classification task for identifying emotion categories. MMER performs multilabel emotion recognition, detecting multiple emotions simultaneously. MABSA involves extracting aspects and opinions to determine sentiment polarity, categorizing it as information extraction. In terms of analysis granularity, MERC and MECPE focus on utterances and speakers within a conversation, while MSA and MMER concentrate on sentence-level information within a document. MABSA, on the other hand, focuses on aspects within comments. Some studies infer fine-grained sentiment from coarse-grained sentiment [209, 319] or integrate tasks of different granularities into a unified training framework [307]. Due to these differences in granularity, the contextual information varies as well. For instance, in MABSA, the context includes the comment along with any associated images and short descriptions of aspects, whereas in MERC, the context encompasses the entire conversation and speaker information. In terms of application scenarios, MSA, MMER, and MABSA are used for public opinion analysis and mining user experiences related to products or services. MERC and MECPE help machines understand and mimic human behaviors, generating empathetic responses in dialogue agents. While many tasks are context-specific, there is a growing trend toward unified frameworks for analyzing human emotions across diverse settings (e.g., task type and emotion granularity) [38, 138].
X Future work
We outline directions for future work in multimodal affective computing from Transfer Learning with Multimodal Pre-trained Model, Unification of Multimodal Affective Computing tasks, Model with External Knowledge Distill, and Affective Computing with Less-studied Modalities.
Unification of Multimodal Affective Computing tasks
Recent advances have made significant strides by unifying related yet distinct tasks into a single framework [26, 212, 215]. For example, T5 [60] integrates various NLP tasks by representing all text-based problems in a text-to-text format, achieving state-of-the-art results across numerous benchmarks. These studies highlight the effectiveness of such unified frameworks in enhancing model performance and generalization [214, 320, 214]. Meanwhile, the progress also suggests the potential for unifying multimodal affective computing tasks across diverse application scenarios. First, unification across different granularity–from fine to coarse—has been employed to train models effectively [209]. Second, an increasing number of pre-trained models now handle language, vision, and audio simultaneously, enabling end-to-end processing across single, dual, and multiple modalities [321]. Third, integrating emotion-cause analysis with emotion recognition in a multimodal setting within a single architecture can enhance their mutual indications and improve overall performance.
Transfer Learning with External Knowledge Distill
In the field of affective computing, incorporating external knowledge such as sentiment lexicons and commonsense knowledge is crucial for a deeper understanding of emotional expressions within the context of social norms and cultural backgrounds [322]. For example, The expression and perception of emotion also varies across cultures, both in text and in face-to-face communication [323]. These differences are critical for cross-cultural sentiment analysis.
Affective Computing with Less-studied Modalities
Natural language (spoken and written), visual data (images and videos), and auditory signals (speech, sound, and music) have long been central to multimodal affective computing. Recently, new sensing data types, like haptic and ECG signals, are gaining attention [324]. Haptic signals, which involve touch and convey sensory and emotional attributes, enhance user experiences in areas such as gaming, virtual reality, and mobile apps [325]. These signals offer immediate feedback and can improve user engagement. As research progresses, less-studied modalities like haptic will likely become crucial, complementing established methods and advancing the field of affective computing.
XI Conclusion
Multimodal Affective Computing has emerged as a crucial research direction in artificial intelligence with significant progress in understanding and interpreting emotions. This survey provides a comprehensive overview of the diverse tasks associated with multimodal affective computing, covering its research background, definitions, related work, technical approaches, benchmark dataset, and evaluation metrics. We group multimodal affective computing across MSA, MERC, MABSA and MMER tasks into four categories: multi-task learning, pre-trained modal, enhanced knowledge and context information. Additionally, we summarize the consistency and differences among various affective computing tasks. Also, we report the inherent challenges in multimodal sentiment analysis and explore potential directions for future research and development.